Need Help in SHM Problem:?

A particle executing shm along y axis such that its velocities are v1 and v2 at distances y1 and y2 from the mean position. Calculate the time period of the motion.

Grade:12

2 Answers

sarim khan
9 years ago
w=ang freq,A=amplitude
v1=w(A^2-y1)^1/2
v2=w(A^2 -y2)^1/2
eliminating A from above equations we get
w^2=(v1^2-v2^2)/(y2^2-y1^2)
From this you can get T=2pi/w
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

w = ang freq,
A=amplitude
v1 = w(A^2-y1)^1/2
v2 = w(A^2 -y2)^1/2
eliminating A from above equations we get
w^2 = (v1^2-v2^2)/(y2^2-y1^2)
From this you can get T = 2pi/w

Thanks and Regards